OpenStudy (anonymous):

write as a single log log(x-1)+log2-(1/2logx)

OpenStudy (anonymous):

\[ \log((x-1)+100-x ^{1/2})\]

OpenStudy (anonymous):

how did you get 100

OpenStudy (anonymous):

oh right base 10^2

OpenStudy (dumbcow):

loga + logb = logab loga - logb = log a/b log a^n = nlog a use these rules to get \[=\log_{}\frac{2(x-1)}{\sqrt{x}} \]

OpenStudy (anonymous):

dumbcow is correct

OpenStudy (anonymous):

sinceits addition you multiply amd subtraction you divide i am wrong

OpenStudy (anonymous):

thats correct thanks guys

OpenStudy (anonymous):

i have another one posted maybe you could help =)